Can someone show me how to do this with a calculator?
Use a calculator to approximate 8√5+8√3 to the nearest hundredth.
8√5+8√3


Answer by checkley75(3666)

8SQRT5+8SQRT3=8*2.236+8*1.732=17.889+13.8564=31.74 ANSWER.